# MASH: Masked Anchored SpHerical Distances for 3D Shape Representation and Generation

### v1

@@ 4.3. Shape Generation
As a compact parametric representation, MASH provides a natural embedding for 3D shapes. We show how MASH can be further used for shape generation in two most common settings, i.e., category-conditioned and image-conditioned. For both settings, we chose the same network backbone as in 3DShape2VecSet (Zhang et al., 2023), as MASH consists of an unordered set of anchor parameters. Then, for the category-conditional model, we create a learnable embedding vector for each category and inject it as a condition into the network. For the image-conditional model, we extract the image features using the pre-trained DINOv2-ViT-B/14 model (Oquab et al., 2023). More details are provided in the supplementary material.
@@
We train the category-conditioned model on the ShapeNet-V2 dataset (Chang et al., 2015) and compare our method with LN3Diff (Lan et al., 2025) and 3DShape2VecSet (S2V) (Zhang et al., 2023). Thanks to the compactness of MASH, the training of our network converges much quicker than other methods. Compared to S2V, both the training and sampling time of our model is less than one-third, with the same network backbone.
To measure the mesh quality of the generative model, we follow CLAY (Zhang et al., 2024) to use Render-FID, Render-KID, P-FID, and P-KID as metrics, computed on 200 generated shapes and shapes random selected from the ShapeNet-V2 dataset for each category. Quantitative and qualitative comparisons are presented in Table 2 and Figure 10, respectively. From the comparisons, we can see that the shapes generated based on MASH exhibit a high diversity and have nice geometric details.
Moreover, as an explicit patch-based representation, one unique feature of MASH is that a subset of its anchors can be fixed during the generation, which enables novel applications like part-conditional generation and shape blending. More specifically, during the training of our category-conditional model, we replace a random proportion of the initial noise fed into the generative model with the parameters of the ground truth MASH in 80% of the steps. With this modification, our trained category-conditioned generative model can naturally support completion and blending tasks. Some visual examples are shown in Figure 7. We can see that the generative model fully utilizes the information of the fixed anchors and attempts to provide reasonable completion or blending results, with all fixed parts are nicely preserved. Note that the slight change of local geometry of the given part is mainly due to surface extraction.
@@
We train the image-conditioned model on the Objaverse-82K dataset (Deitke et al., 2023) and compare our method to baselines including InstantMesh (Xu et al., 2024), Make-A-Shape (Hui et al., 2024) and Hunyuan3D-1 (Yang et al., 2024). To perform a quantitative comparison, we randomly select 100 3D shapes from the Objaverse dataset and render an image from a random viewpoint for each shape to serve as the input of different methods. Then, we normalize the 3D shapes generated by different methods together with the 3D shapes in the dataset and manually align them to eliminate any possible orientation inconsistencies between the generated results and the ground truth. Other than the metrics used for category-conditioned generations, we also introduce Chamfer Distance (CD), Earth Mover’s Distance (EMD), F-Score, and ULIP-I as additional metrics. The quantitative and qualitative comparisons are shown in Table 3 and Figure 11, respectively. Our method obtains generally better results. Both InstantMesh and Hunyuan3D-1 take the input image to generate multi-view images first and then reconstruct the 3D shape from those images, thus they highly depend on the geometric consistency across multiple views. Compared to other methods, the 3D shapes generated based on MASH can better preserve the topology of the object in the image and better recover the full geometric shape.
[…]
@@ 5. Conclusion
As a preliminary attempt, our representation and corresponding optimization still have several limitations. Our current MASH representation only considers geometry and it would be interesting to explore ways to further incorporate textures. For the optimization, the current initialization evenly distributes anchors over the given data, which sometimes lead to sub-optimal fitting results. Adapting the anchor distribution to the given shape will be a promising direction for future work. Additionally, our current optimization strategy is relatively sensitive to outliers, which are commonly presented in range sensor scans and can cause reconstruction artifacts. Enhancing the robustness and efficiency of optimization warrants further in-depth investigation.

### v2

@@ 4.3. Shape Generation
Since MASH is discriminative yet compact, it can naturally serve as embeddings for 3D shapes and has the potential to make learning and inference more efficient compared to implicit representations. We show how MASH can be further used for shape generation in two most common settings, i.e., category-conditioned and image-conditioned. Considering that MASH consists of a set of unordered surface patches, each represented by a fixed-length set of parameters, making it fit the architecture of 3DShape2VecSet (Zhang et al., 2023). Therefore we also use the same conditioning for shape generation as in 3DShape2VecSet. Specifically, for the category-conditional model, we create a learnable embedding vector for each category and inject it as a condition into the network. For the image-conditional model, we extract the image features using the pre-trained DINOv2-ViT-B/14 model (Oquab et al., 2023). To further improve the parameter distribution for generation, we applied an invertible piecewise linear transform to convert it into normal distributions, which is a modified version of the PowerTransformer in sklearn (Buitinck et al., 2013) to ensure that the transformation error is below 1e-6 for better precision. More details are provided in the Supplementary.
@@
Our category-conditioned model is trained on ShapeNet-V2 (Chang et al., 2015). We compare our method with LN3Diff (Lan et al., 2025) and 3DShape2VecSet (S2V) (Zhang et al., 2023). Thanks to the compactness of MASH, the training of our network converges much quicker than other methods. Compared to S2V, both the training and sampling time of our model is less than one-third, with the same network backbone.
Following CLAY (Zhang et al., 2024), we measure quality of the generative meshes using Render-FID, Render-KID, P-FID, and P-KID, computed on 200 generated shapes and shapes randomly selected from ShapeNet-V2 for each category. Quantitative and qualitative comparisons are presented in Table 2 and Figure 10, respectively. The results clearly show that shapes generated via MASH exhibit greater diversity with better geometric details.
Moreover, as an explicit patch-based representation, one unique feature of MASH is that a subset of its anchors can be fixed during the generation, which enables novel applications like part-conditional generation and shape blending. More specifically, during the training of our category-conditional model, we replace a random proportion of the initial noise fed into the generative model with the parameters of the ground truth MASH in 80% of the steps. With this modification, our trained category-conditioned generative model can naturally support completion and blending tasks. Some visual examples are shown in Figure 7, where we use the orange color to highlight the retained shape parts, and cyan to represent the entire generated 3D shapes. We can see that the generative model fully utilizes the information of the fixed anchors and attempts to provide reasonable completion or blending results, with all fixed parts are nicely preserved. Note that the slight change of local geometry of the given part is mainly due to surface extraction.
@@
We train the image-conditioned model on the Objaverse-82K dataset (Deitke et al., 2023) and compare our method to baselines including InstantMesh (Xu et al., 2024), Make-A-Shape (Hui et al., 2024) and Hunyuan3D-1 (Yang et al., 2024). To perform a quantitative comparison, we randomly select 100 3D shapes from the Objaverse dataset and render an image from a random viewpoint for each shape to serve as the input of different methods. Then, we normalize the 3D shapes generated by different methods together with the 3D shapes in the dataset and manually align them to eliminate any possible orientation inconsistencies between the generated results and the ground truth. Other than the metrics used for category-conditioned generations, we also introduce Chamfer Distance (L1-CD) and ULIP-I as additional metrics. The quantitative and qualitative comparisons are shown in Table 3 and Figure 11, respectively. Our method obtains generally better results. Both InstantMesh and Hunyuan3D-1 take the input image to generate multi-view images first and then reconstruct the 3D shape from those images, thus they highly depend on the geometric consistency across multiple views. Compared to other methods, the 3D shapes generated based on MASH can better preserve the topology of the object in the image and better recover the full geometric shape.
@@ 5. Conclusion
As a preliminary attempt, our representation and corresponding optimization still have several limitations. Our current MASH representation only considers geometry and it would be interesting to explore ways to further incorporate textures. For the optimization, the current initialization evenly distributes anchors over the given data, which sometimes lead to sub-optimal fitting results. Adapting the anchor distribution to the given shape will be a promising direction for future work. At last, since the current datasets employed for training our methods and the scale of the generative network are relatively small, the geometric details of the generated results still leave much room for improvements. Training a more powerful generative model on a larger-scale dataset is a worthwhile pursuit for future work.
Acknowledgements.
We thank the anonymous reviewers for their valuable comments. This work is supported by the National Key R&D Program of China (2024YFB2809102), National Natural Science Foundation of China (U24B20154, 62025207, 62322207), Joint NCSF-ISF Research Grant (3077/23), and Shenzhen University Natural Sciences 2035 Program (2022C007).